OpenStudy (anonymous):

how do i solve 1-cos(theta)=1/2?

OpenStudy (anonymous):

is 1-cos = 1*(-cos)?

OpenStudy (anonymous):

1-cos(theta)=1/2 -cos(theta)=1/2+1 cos(theta)=-(1/2+1) theta=arccos(-(1/2+1))